### v1

# Holodeck 2.0: Vision‑Language‑Guided 3D World Generation with Editing

Abstract
3D scene generation plays a crucial role in gaming, artistic creation, virtual reality and many other domains. However, current 3D scene design still relies heavily on extensive manual effort from creators, and existing automated methods struggle to generate open-domain scenes or support flexible editing. As a result, generating 3D worlds directly from text has garnered increasing attention. In this paper, we introduce Holodeck 2.0, an advanced vision-language-guided framework for 3D world generation with support for interactive scene editing based on human feedback. Holodeck 2.0 can generate diverse and stylistically rich 3D scenes (e.g., realistic, cartoon, anime, and cyberpunk styles) that exhibit high semantic fidelity to fine-grained input descriptions, suitable for both indoor and open-domain environments. Holodeck 2.0 leverages vision-language models (VLMs) to identify and parse the objects required in a scene and generates corresponding high-quality assets via state-of-the-art 3D generative models. It then iteratively applies spatial constraints derived from the VLMs to achieve semantically coherent and physically plausible layouts. Human evaluations and CLIP-based assessments demonstrate that Holodeck 2.0 effectively generates high-quality scenes closely aligned with detailed textual descriptions, consistently outperforming baselines across indoor and open-domain scenarios. Additionally, we provide editing capabilities that flexibly adapt to human feedback, supporting layout refinement and style-consistent object edits. Finally, we present a practical application of Holodeck 2.0 in procedural game modeling, generating visually rich and immersive environments, potentially boosting efficiency.
Figure 1: Examples of stylistically varied 3D scenes generated by Holodeck 2.0. Captions are shortened versions of the original inputs (50–300 words). Holodeck 2.0 faithfully captures fine-grained details from textual descriptions.
Introduction
The ability to generate complex 3D environments from natural language descriptions represents a fundamental challenge at the intersection of computer vision, natural language processing, and computer graphics. As virtual worlds become increasingly central to gaming, film production, architectural visualization, and training embodied AI agents, the demand for efficient, high-quality 3D content creation has never been greater. However, the current practice of 3D scene design still relies predominantly on extensive manual effort by creators. Therefore, the task of generating 3D worlds directly from text or images has attracted widespread attention.
Some current approaches focus on directly generating 3D scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images (Poole et al. 2022; Lin et al. 2023; Chung et al. 2023). However, such methods generate scenes that lack separable and manipulable objects, making them difficult to edit directly and limiting their applicability in game design, embodied AI training, etc. Another class of methods focuses on selecting suitable 3D objects from asset datasets and then assembling them into a 3D scene (Paschalidou et al. 2021; Yang et al. 2024; Hu et al. 2024; Aguina-Kang et al. 2024). For example, Holodeck (Yang et al. 2024) uses Large Language Models (LLMs) to parse from the input text the objects required for a scene, then retrieves 3D assets from datasets such as Objaverse (Deitke et al. 2023b) based on Contrastive Language–Image Pre-training (CLIP) (Radford et al. 2021) scores, and finally arranges them into a layout.
Specifically, existing approaches face three critical limitations: (1) asset-based methods are constrained by the quality and diversity of available 3D model libraries, with even large datasets like Objaverse-XL (Deitke et al. 2023a) suffering from inconsistent quality and missing textures; (2) current systems struggle with fine-grained style control, unable to maintain consistent artistic styles (e.g., anime, cyberpunk) across all objects in a scene; and (3) most methods focus exclusively on indoor environments, lacking the flexibility to generate open-world scenarios like fantasy landscapes or futuristic cities.
Recent breakthroughs in vision-language models (VLMs) and 3D generative models present an unprecedented opportunity to address these limitations. VLMs now possess sophisticated spatial reasoning capabilities that can parse complex scene descriptions and generate physically plausible object arrangements. Simultaneously, state-of-the-art 3D generation models like Hunyuan3D can create high-quality, stylistically consistent assets from text or image prompts (Hunyuan3D et al. 2025; Lai et al. 2025). However, these advances have not yet been effectively combined into a unified scene generation framework.
Facilitated by 3D asset generation models and the inference capabilities of VLMs, Holodeck 2.0 can produce high-quality 3D scenes with internal stylistic coherence for both indoor settings (e.g., music studios, playrooms) and open domains (e.g., airship docks, courtyards), as shown in Figure 1. Meanwhile, Holodeck 2.0 can generate the same type of 3D scenes in different styles, such as realistic, cartoon, and steampunk. The objects in Holodeck 2.0 scenes exhibit high quality and closely match their descriptions, for example, “A cartoon-style snail-shaped snack booth offers glowing drinks.” In terms of spatial reasoning, Holodeck 2.0 fully leverages the spatial understanding capabilities of VLMs, iteratively generating and applying spatial constraints to produce layouts that match both textual and visual descriptions.
Beyond research contributions, we demonstrate Holodeck 2.0’s practical applicability through direct integration with commercial game engines. We showcase how a complex museum scene complete with artifacts like the Temple of Dendur, Venus de Milo, and The Thinker—can be generated from detailed text descriptions and seamlessly imported into Unreal Engine with full interactivity. This demonstrates how our system can accelerate 3D content creation workflows for game development and other applications. We demonstrate Holodeck 2.0’s effectiveness through large-scale human evaluations and CLIP-based automated scoring. Across 60 test cases spanning both indoor and open-domain scenes, Holodeck 2.0 consistently outperforms the baseline Holodeck in both human evaluations and CLIP-based automated scoring, with its advantage being particularly pronounced in open-domain scenes.
This paper makes the following contributions:
  1. 1.
    We present a unified framework that uniquely combines VLM-guided scene decomposition with generative 3D asset creation, using VLMs across multiple stages—from visual reference generation to object parsing, spatial constraint generation, and iterative layout refinement.
  2. 2.
    We demonstrate state-of-the-art performance on both indoor and open-domain scene generation with consistent artistic styles (realistic, cartoon, anime, cyberpunk) through extensive human and automated evaluations.
  3. 3.
    We introduce an interactive editing system that leverages generative models for style-consistent object replacement and addition, beyond traditional layout adjustments.
  4. 4.
    We validate practical applicability through integration with commercial game engines and demonstrate significant quality improvements: 110% for indoor scenes and 139% for open-domain scenes compared to prior work.
Our code and generated scenes will be made publicly available to facilitate future research.
Figure 2: Overview of Holodeck 2.0. Given a text input, Holodeck 2.0 generates 3D scenes via three modules, with an optional editing module for interactive updates based on human language feedback.
Related Work
Language-Guided 3D Scene Generation
Early work in this area is primarily symbolic and rule‑based, requiring extensive engineering effort and exhibiting very limited generalization ability, for example, Wordseye (Coyne and Sproat 2001) and other systems (Chang, Savva, and Manning 2014; Ma et al. 2018). With the emergence of LLMs and VLMs, open‑vocabulary 3D scene generation has become possible, greatly enhancing the flexibility of scene creation. Holodeck (Yang et al. 2024) employs LLMs to get the object descriptions that should appear in a scene based on the input text, and to generate spatial constraints among those objects. SceneCraft (Hu et al. 2024) also leverages LLMs and VLMs in a similar way, but it generates Blender code and iteratively optimizes it to produce 3D scenes. However, both Holodeck and SceneCraft rely on CLIP (Radford et al. 2021) to retrieve 3D objects from asset libraries such as Objaverse (Deitke et al. 2023b) and TurboSquid (TurboSquid 2024), and the variety and quality of these assets somewhat limit their capability to generate open‑world 3D scenes. Currently, larger or better‑curated 3D asset libraries have emerged (Deitke et al. 2023a; Lin et al. 2025). However, they still suffer from overall low asset quality, missing textures, and limited variety and styles of objects. For 3D scene spatial layout, LayoutGPT (Feng et al. 2023) is a method that directly generates 3D object layout relationships via LLM prompting, but often neglects physical plausibility. LayoutVLM (Sun et al. 2025) is another method that uses VLMs to generate 3D layouts by formulating objective functions that can be differentiably optimized. For 3D scene editing, recent works such as BlenderGPT (BlenderGPT 2023) and BlenderAlchemy (Huang, Yang, and Guibas 2024) have applied LLMs or VLMs to editing, but these models primarily focus on material replacement and perform poorly at placement or layout.
Text-to-3D Diffusion Models
The emergence of diffusion models (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Peebles and Xie 2023; Rombach et al. 2022) marks a significant advancement in the field of generative modeling. Current methods leverage diffusion models to directly generate 3D objects or scenes represented by Neural Radiance Fields (NeRFs) or Gaussian splats from text or images, such as DreamFusion (Poole et al. 2022) and Magic3D (Lin et al. 2023). However, such methods generate scenes that lack discrete, manipulable objects, making them difficult to edit directly and unsuitable for applications like game design or embodied AI training. When focusing solely on the generation of individual 3D assets, diffusion models offer significant advantages. Recent methods have achieved breakthroughs in generating high‑quality 3D assets (Xiang et al. 2025; Zhang et al. 2023, 2024; Zhao et al. 2023; Li et al. 2024), with Hunyuan3D (Hunyuan3D et al. 2025; Lai et al. 2025) being a typical example. It can generate high‑quality, mesh‑based individual 3D assets from text or images.
Figure 3: Examples of the Scene Analysis Module and the Object Generation Module. Holodeck 2.0 can generate customized, stylistically diverse 3D objects that precisely match fine-grained textual descriptions.
Holodeck 2.0
Holodeck 2.0 is a vision-language-guided system, facilitated by VLMs and 3D generative models, that can produce and edit diverse, customizable, and stylistically varied open-domain 3D scenes in Blender.
As shown in Figure 2, Holodeck 2.0 systematically generates open-domain 3D scenes from text or images utilizing the following specialized modules: (1) Scene Analysis Module parses the object properties and background information that should appear in the scene from input text or images; (2) Object Generation Module leverages 3D asset generation models to produce 3D assets from quality‑controlled object images; (3) Scene Generation Module iteratively generates semantically and physically coherent object layouts based on spatial constraints; (4) Scene Editing Module automatically applies personalized adjustments to the scene based on human feedback.
In the following, we describe how each module of Holodeck 2.0 operates, specifically illustrating how it coordinates VLMs and 3D asset generation models to accomplish scene generation and editing. We also present its direct application to game design. Detailed information about Holodeck 2.0 is also available in the supplement.
Figure 4: Example outputs of the Scene Generation Module. Holodeck 2.0 can generate appropriate spatial constraints and optimal positioning for object layouts.
Scene Analysis
The Scene Analysis Module, shown in the first panel of Figure 2, is designed to parse object properties that should appear in the scene from text or images.
Text to Scene Image. Given a text input, the Scene Analysis Module uses VLMs to generate a corresponding reference image. This image generation process is vital for the subsequent creation of uniform‑style assets. Moreover, visual inputs can substantially enhance the spatial awareness of large reasoning models. In Figure 3, we present several examples of scene reference images generated from text.
Scene Image to Object Images. After obtaining the scene image, the Scene Analysis Module employs GPT-o3 (OpenAI 2025) to infer object properties (name, position, rotation, size, visual description, etc.) from the text input and its corresponding scene reference image. Then we prompt the VLMs to generate a clear, transparent frontal image of each object appearing in the scene image. This avoids the issues common to current image segmentation algorithms (Kirillov et al. 2023; Ren et al. 2024), such as inaccurate segmentation, blurry edges, and difficulty handling occluded objects. After generating individual object images, we use GPT-o3 to perform quality control, removing smaller-object images redundantly embedded in larger-object images (e.g., removing the separate faucet when it is already present in the bathtub image). In practice, this approach generates accurate, high‑quality object images from the reference image as shown in Figure 3. Subsequently, we generate the scene’s background image based on the reference image, focusing primarily on the ground texture.
Object Generation
The Object Generation Module, shown in the second panel of Figure 2, utilizes 3D asset generation models to generate 3D assets from quality-controlled object images. For 3D assets generation, image-to-3D is faster and more stable than direct text-to-3D. After obtaining the object images from the Scene Analysis Module, we use Hunyuan3D 2.1 to generate the corresponding 3D asset for each object. It is one of the most powerful open‑source 3D generation models that supports local deployment. In practice, compared to searching for 3D objects in Objaverse, generative models can produce higher‑quality 3D assets that more faithfully match the descriptions (Figure 3), for example, “a plush, L-shaped sectional sofa adorned with colorful throw pillows” and “a cartoon-style donut-shaped bench”.
Figure 5: The Scene Editing Module of Holodeck 2.0 can add/delete objects, change layouts, and change object styles based on human language feedback.
Figure 6: 3D Scenes generated by Holodeck 2.0 can be seamlessly integrated into standard game production pipelines. From left to right: the Holodeck 2.0 output—a museum hall featuring the Temple of Dendur, the Venus de Milo, The Thinker, etc.; the same scene imported into Unreal Engine and developed into a museum-guide game; and the in-game interface showcasing high-fidelity, fully interactive assets.
Scene Generation
The Scene Generation Module, shown in the third panel of Figure 2, can generate semantically and physically coherent object layouts through iterative cycles between generated spatial constraints and a Depth‑First-Search (DFS) solver.
Spatial Constraint Generation. The Scene Generation Module first employs GPT‑o3 to produce a series of spatial constraints between objects based on the text input, the scene reference image, and object properties. We define a total of ten spatial relations, which are divided into the following four categories: (1) Relative: in front of, behind, left of, right of, side of; (2) Distance: near, far; (3) Vertical: on, above and (4) Rotation: face to. As shown in Figure 4, large reasoning models can generate realistic and spatially plausible constraints for objects.
Layout Generation. For layout generation, we employ a DFS solver to compute object placements based on object properties and spatial constraints between objects. We define mathematical formulations to enforce each spatial constraint (for example, “face to” means that the angle between the source object’s orientation and the line connecting the centers of the two objects is less than 10 degrees). Objects should satisfy these spatial constraints while avoiding physical collisions. To fully leverage GPT-o3’s spatial reasoning capabilities, we prioritize the initial object positions inferred by the model. It is worth noting that, when the DFS solver encounters spatial constraints that are difficult to satisfy simultaneously, the Scene Generation Module re‐prompts GPT‑o3 using the DFS solver’s feedback to update the objects’ spatial constraints. The DFS solver and the reasoning model then form an iterative loop until a valid layout is found. This approach significantly enhances Holodeck 2.0’s layout capabilities. We present several examples of position plots for the bounding boxes produced by the solver (Figure 4).
Scene Editing
The Scene Editing Module, shown as a submodule in the third panel of Figure 2, is designed to automatically apply personalized adjustments to the scene based on user language feedback when needed. It can operate at both the scene layout level and the 3D asset level (Figure 5).
Scene Layout Level. If users wish to personalize the generated scene layout after it has been created, they can provide a natural-language edit instruction. Then, it is incorporated into the original workflow to regenerate the spatial constraints between objects, and the DFS solver iteratively computes the updated layout.
3D Asset level. Utilizing 3D generative models, we support asset‑level editing. Because our scenes are built by assembling individual 3D assets, we can easily delete objects in response to user language feedback. Users can also replace objects, for example, “Change the torii gate to a pastel gradient color scheme and hang decorative bells on it.” The Scene Editing Module then uses that language feedback to regenerate a style‑appropriate 3D asset for the specified object and substitute it into the scene. When adding new objects, the module regenerates suitable 3D assets and recalculates spatial constraints to maintain coherent scene layouts.
Applications in Gaming
We explore the potential of directly integrating Holodeck 2.0-generated scenes into commercial game engines for game development. To demonstrate this, we use Holodeck 2.0 to construct a detailed museum exhibition hall based on rich textual descriptions specifying each artifact’s type, material, style, and placement. Holodeck 2.0 produces a 3D museum scene that closely matches the text, featuring the Temple of Dendur at its center, surrounded by over a dozen objects such as the Venus de Milo and a Shang‑dynasty bronze ding cauldron. This generated 3D scene can then be directly imported into professional game engines such as Unreal Engine, enabling quick and intuitive development of interactions between the game character and artifacts. Such a virtual museum game can serve both educational and entertainment purposes. The results in Figure 6 demonstrate Holodeck 2.0’s high-quality, highly faithful scene generation from detailed text, highlighting its potential for direct application in game content creation.
Evaluation
To systematically evaluate the quality of 3D scenes generated by Holodeck 2.0, we employ a combination of human evaluations (involving more than 300 responses, the majority of which are from undergraduate and graduate students) and CLIP‑based automated scoring. Holodeck (Yang et al. 2024) is a representative method for 3D scene generation that retrieves objects from 3D asset libraries and is one of the most commonly used baselines in recent work. Although it primarily focuses on indoor scene generation, its open‑vocabulary text input makes it applicable to open‑domain scenes to some extent. Therefore, we choose to directly compare with the Holodeck method. Specifically, we generated 60 test cases spanning both indoor and open‑domain scenes. Our comparisons show that Holodeck 2.0 outperforms the baseline Holodeck in both indoor and open‑domain settings, with the advantage being more pronounced in open‑domain scenarios. In addition, we perform ablation studies on the Scene Generation Module. We further compare the Scene Editing Module with the baseline method, BlenderAlchemy (Huang, Yang, and Guibas 2024). The detailed results are shown below.
Comparative Analysis
Setup. We use LLMs to generate 30 input texts as test cases for indoor scenes and another 30 for open-domain scenes (see the supplement for examples of detailed descriptions, each averaging about 50 words). These indoor scenes span different categories, including bedrooms, gyms, playrooms, museums, and laboratories. These open-domain scenes are varied, including architectural complexes, natural landscapes, and futuristic environments. They also exhibit different styles, such as realistic, cartoon and steampunk. We then feed each text into both Holodeck and Holodeck 2.0 to produce 60 paired scenes (30 paired indoor scenes and 30 paired open-domain scenes) for human evaluation and CLIP‑based scoring. In the human evaluation experiment, we render each 3D scene from the least occluded viewpoint using identical parameters, then present the rendered images to annotators in shuffled order within each pair. Annotators score each pair on a scale from 1 to 10 along two dimensions: (1) Asset level: the overall quality of the 3D assets displayed in the scene; (2) Scene level: the overall quality of the scene with respect to the input description. For CLIP-based scoring, we measure the embedding similarity (CLIP ViT-B/32) between each generated scene’s rendered image and the input text to evaluate the quality of the generation.
Methods Asset-level Scene-level CLIP Score
Holodeck 4.28 3.67 0.253
Holodeck 2.0 8.00 7.71 0.299
Table 1: Evaluation results on Indoor Scenes. We have human-evaluated asset-level and scene-level scores, as well as the automatically computed CLIP Score.
Methods Asset-level Scene-level CLIP Score
Holodeck 3.77 3.06 0.256
Holodeck 2.0 7.58 7.32 0.307
Table 2: Evaluation results on Open-domain Scenes.
Results. Each of the indoor and open-domain scene evaluation surveys is completed by 105 annotators. Table 1 and Table 2 present the quantitative scores (human evaluation and CLIP-based scoring) for indoor scenes and open-domain scenes, comparing Holodeck 2.0 and Holodeck. Figure 7 illustrates the full 1–10 rating distributions from the human evaluation, showing how scores are spread across both asset and scene levels. Figure 8 shows the CLIP score distributions. These results show Holodeck 2.0 outperforms Holodeck in both indoor and open‑domain settings, with the advantage being more significant in the latter, demonstrating its strong capability in scene generation tasks.
Figure 7: Human evaluation on 1–10 rating distributions at asset and scene levels for indoor (left) and open-domain (right) scenes, comparing baseline Holodeck (blue hues) versus Holodeck 2.0 (red hues).
Figure 8: Comparison of CLIP score distributions for Holodeck (baseline) versus Holodeck 2.0 (ours). Kernel density estimates are shown for 30 indoor scenes (left) and 30 open-domain scenes (right), with dashed lines marking the mean score of each method.
Figure 9: Comparative human evaluation of Holodeck 2.0 across two experiments. Pie charts display annotator preference percentages and counts for Ablation Study (left) and Scene Editing Comparison (right).
Ablation Study on the Scene Generation Module
Setup. We also conduct an ablation study on Holodeck 2.0’s Scene Generation Module. We evaluate two configurations: (1) Using only the initial layouts from the VLM; (2) Using the full Scene Generation Module. For each input text from the Comparative Analysis, we generate 2 scenes (with the same 3D assets) under both settings and then assess their quality using human evaluation. We asked human evaluators to choose among 60 pairs of scene layouts, with three options available: “Prefer A”, “Prefer B”, and “Tie”.
Results. 94 annotators complete the survey, with each annotator answering all 60 questions. Human evaluation shows that the full Scene Generation Module is indispensable for effective 3D scene generation. As the left panel of Figure 9 shows, 67.93% of the choices favor the scenes after our layout, 12.89% prefer the initial placements generated directly by the VLM, and the remaining 19.18% are ties.
Evaluation on Scene Editing
Previous 3D scene editing methods, such as BlenderGPT and BlenderAlchemy (Huang, Yang, and Guibas 2024), have mainly focused on material design and lighting systems, but struggle with style replacement and object placement. By fully utilizing 3D asset generation models, we accomplish what methods like BlenderAlchemy struggle to achieve, such as adding objects in specific artistic styles or replacing existing objects with customized stylistic variations.
For object placement editing, since BlenderAlchemy can achieve this by editing Blender scripts with VLMs, we conduct a human evaluation to compare Holodeck 2.0 against BlenderAlchemy in placement editing tasks. In our control experiment: (1) We perform 10 representative placement editing tests for both BlenderAlchemy and Holodeck 2.0. (2) Annotators are then required to choose their preferred version between the two edited scenes.
197 annotators complete the survey, with each annotator answering all 10 questions. As the right panel of Figure 9 shows, 71.12% of annotators prefer our edited scenes, while only 28.88% favor BlenderAlchemy’s outputs. This demonstrates that our integrated approach combining DFS and VLM-generated spatial constraints for layout editing outperforms the method relying solely on VLM-based editing and evaluation.
Conclusions and Discussions
We propose Holodeck 2.0, a vision-language-guided system, equipped with 3D asset generation models, that can generate and edit diverse, customizable, and stylistically varied 3D worlds. We validate the effectiveness of Holodeck 2.0 through human evaluations and CLIP-based scoring. We also show its potential in game modeling. In future work, we intend to improve Holodeck 2.0’s capacity for generating more complex and finely detailed 3D worlds, thereby facilitating the direct creation of large-scale game environments.
Acknowledgements
We thank Ben Swanson for his help in improving our understanding of 3D generative models and game engines, and for providing helpful suggestions on the manuscript.
Appendix
Details of Holodeck 2.0
Details of the Scene Analysis Module
In the Scene Analysis Module, Holodeck 2.0 first uses GPT-Image-1 to generate a reference image based on the input text. The prompt is shown below.
Listing 1 Prompt - Generate Reference Image
1prompt = (
2 f"{description}. Render in {style} style. 3-D view: x->right, y->backward, z->up. "
3 "Well-lit, no extra objects."
4)
Then, Holodeck 2.0 feeds the input text and the reference image into GPT-o3, which outputs a JSON file containing object properties. The prompt is shown below. We also illustrate the structure of the JSON file through an example.
Listing 2 Prompt - Generate Object Properties
1class Vec3(BaseModel):
2 x: float
3 y: float
4 z: float
5
6class SceneItem(BaseModel):
7 id: str
8 name: str
9 position: Vec3
10 rotation: Vec3
11 size: Vec3
12 visual_description: str
13
14class SceneData(BaseModel):
15 items: List[SceneItem]
16
17SCHEMA_JSON = SceneData.model_json_schema()
18
19
20INSTRUCTION = (
[…]
31 "Each entry in the items list must correspond to a complete object (e.g. a building, a sculpture)--"
[…]
34)
Listing 3 JSON File — Object Properties
1{
2 "items": [
3 {
4 "id": "bed1",
5 "name": "King Bed",
6 "position": {
7 "x": 0.0,
8 "y": 1.0,
9 "z": 0.4
10 },
11 "rotation": {
12 "x": 0.0,
13 "y": 0.0,
14 "z": 0.0
15 },
16 "size": {
17 "x": 1.92,
18 "y": 1.94,
19 "z": 1.2
20 },
21 "visual_description": "A king-sized bed with a light-beige upholstered headboard, dressed in a white duvet and multiple decorative pillows."
22 },
23 {
24 "id": "nightstand_right",
25 "name": "Right Nightstand",
26 "position": {
27 "x": 1.2,
28 "y": 1.0,
29 "z": 0.35
30 },
31 "rotation": {
32 "x": 0.0,
33 "y": 0.0,
34 "z": 0.0
35 },
36 "size": {
37 "x": 0.52,
38 "y": 0.95,
39 "z": 0.6
40 },
41 "visual_description": "A matching light-wood minimalist nightstand with one drawer and open lower shelf."
42 },
43 {
44 "id": "lamp_left",
45 "name": "Left Table Lamp",
46 "position": {
47 "x": -1.2,
48 "y": 1.0,
49 "z": 0.9
50 },
51 "rotation": {
52 "x": 0.0,
53 "y": 0.0,
54 "z": 0.0
55 },
56 "size": {
57 "x": 0.34,
58 "y": 0.34,
59 "z": 0.6
60 },
61 "visual_description": "A sleek glass-base table lamp with a white drum shade."
62 },
63 ...
64 ]
65}
Then, Holodeck 2.0 extracts object names from the JSON file to form a list. This list, along with the scene reference image, is then passed to GPT-Image-1 to generate frontal images for individual objects. The prompt is shown below.
Listing 4 Prompt - Generate Individual Images
1prompt = (
2 f"Please generate ONE PNG image of an isolated front-view {obj_name} "
3 f"with a transparent background, in {style} style, with shapes and style similar to the reference scene. "
4)
After generating the individual images for different objects, Holodeck 2.0 uses GPT-o3 again for quality control. For example, sometimes a generated image of a large object may already include a smaller object (e.g., a bathtub image containing a faucet), while a separate image of the smaller object is also generated. Holodeck 2.0 prompts GPT-o3 to correct such cases by removing redundant images of the smaller objects. The prompt is shown below.
Listing 5 Prompt - Remove Redundant Images
1system_msg = {
2 "role": "system",
3 "content": "You are a helpful assistant that identifies redundant sub-components in a set of images."
4}
5
6class ImageList(BaseModel):
7 filenames: list[str]
8SCHEMA_JSON = ImageList.model_json_schema()
9
10instruction = (
11 f"""
12 I have the following list of image filenames:
13 {json.dumps(png_files, indent=2)}
14 """
15 "Here are all images in the current scene. Some represent smaller parts already "
16 "contained within larger assemblies. Return a JSON list of filenames (with ".png") "
[…]
19)
To better capture background information, Holodeck 2.0 feeds the scene reference image into GPT-Image-1 to extract a background reference image, with a primary focus on the ground background. We aim for the generated images to have a uniform ground surface. The prompt is shown below.
Listing 6 Prompt - Generate Background Image
1prompt = (
2 "Replace the entire image with ONE seamless, tileable PNG of the main floor for indoor scenes, and the main ground for outdoor scenes. "
3 "Using the material and pattern seen in the input photo. Ignore walls, ceiling and decorations."
4 "The texture must be homogeneous, repeating smoothly, and produced at a scale large enough to cover an expansive floor area. Do not add transparency."
5)
Details of the Object Generation Module
In the Object Generation Module, we run the deployed Hunyuan3D 2.1 model locally. Given the individual image of each object, Hunyuan3D 2.1 generates a corresponding 3D asset in GLB-format. Since generating each 3D asset takes approximately 3–5 minutes on an A6000 GPU, Holodeck 2.0 parallelizes the asset generation process based on available resources. After generating the 3D asset for each object, we write the asset’s file path back into the JSON file containing the object properties. We also rescale each 3D asset based on its actual x, y, and z dimensions, using the original z size as a reference. The updated x and y sizes after scaling are then written back into the JSON file.
Details of the Scene Generation Module
In the Scene Generation Module, Holodeck 2.0 first embeds the scene input text along with each object’s ID, name, and visual description (extracted from the JSON file) into the prompt. This prompt, together with the scene reference image, is then fed into GPT-o3 to generate the necessary spatial constraints between objects in the scene. Examples of such spatial constraints are provided below.
Listing 7 JSON File - Spatial Constraints
1[
2 {
3 "type": "relative",
4 "relation": "left of",
5 "source": "nightstand_left",
6 "target": "bed1"
7 },
8 {
9 "type": "relative",
10 "relation": "right of",
11 "source": "nightstand_right",
12 "target": "bed1"
13 },
14 {
15 "type": "relative",
16 "relation": "right of",
17 "source": "dresser1",
18 "target": "bed1"
19 },
20 {
21 "type": "relative",
22 "relation": "on",
23 "source": "lamp_left",
24 "target": "nightstand_left"
25 },
26 {
27 "type": "distance",
28 "relation": "near",
29 "source": "lamp_left",
30 "target": "bed1"
31 },
32 {
33 "type": "relative",
34 "relation": "on",
35 "source": "lamp_right",
36 "target": "nightstand_right"
37 },
38 {
39 "type": "distance",
40 "relation": "near",
41 "source": "lamp_right",
42 "target": "bed1"
43 },
44 {
45 "type": "relative",
46 "relation": "on",
47 "source": "photo_frames1",
48 "target": "dresser1"
49 }
50]
It is worth noting that, to facilitate the operation of our subsequent DFS solver, we specify to GPT-o3 that once an object has been used as a target, it can no longer be used as a source in any following constraints. We also require that constraints with the same source must appear consecutively. The prompt is shown below.
Listing 8 Prompt - Generate Spatial Constraints
1system_prompt = (
2 "You are a spatial relationship analyzer. Generate valid JSON constraints based on scene descriptions, object information and the reference image."
3 "Output the constraints in a strict sequence: once an object has appeared as a target, it must not later appear as a source."
4 "For each object, output all of its source-type constraints in true consecutive order in the list - they must be physically adjacent without any other constraints interleaved.")
5
6user_prompt = f"""
7You are a spatial relationship analyzer. Given a scene description and a list of objects with their IDs, generate spatial constraints between the objects.
8
9Scene Description:
10{description}
11
12Available Objects:
13{objects_text}
14
15Generate spatial constraints in the following JSON format:
16[
17 {{
18 "type": "relative",
19 "relation": "right of|left of|in front of|behind|side of|on|above",
20 "source": "object_id",
21 "target": "object_id"
22 }},
23 {{
24 "type": "distance",
25 "relation": "near|far",
26 "source": "object_id",
27 "target": "object_id"
28 }},
29 {{
30 "type": "rotation",
31 "relation": "face to",
32 "source": "object_id",
33 "target": "object_id"
34 }}
35]
36
37Guidelines:
381. Only include meaningful spatial relationships that match the description
392. Use the exact object IDs from the available objects list
403. Focus on the most important spatial relationships
414. Avoid redundant relationships
425. Return only valid JSON, no additional text
436. For example, if the scene description mentions "a chair to the right of a table", the relation should be right of, with the chair as the source and the table as the target.
447. An object should either choose to be on the ground(default, no need to specify as a constraint) or on/above another object.
458. Use "on" when one object is physically resting on the surface of another object. Use "above" only when the object is floating or suspended in the air above another object, not touching it. If an object is on the ground, do not specify any "on" or "above" relation when it is the source.
46
47Generate the constraints:
48"""
Next, we input the spatial constraints along with the JSON file containing object properties into the DFS solver. The DFS solver works by placing objects one by one based on the dependency structure defined by the constraints. Objects that do not depend on any others are placed first, preferably at their initial positions inferred by the large reasoning model GPT-o3. Subsequently, other objects are placed in order of their dependencies — each object is only placed once all its constraints are satisfied and no collisions are detected. The solver proceeds with this search process until all objects are successfully placed, at which point it outputs the final layout.
If some objects cannot be placed due to conflicting constraints, the DFS solver will report which object’s constraints could not be fully satisfied during the search. We then use this information as part of the prompt (edit_instructions) to GPT-o3 to regenerate the spatial constraints, allowing us to iteratively refine the layout. The prompt used for regenerating the constraints is shown below.
Listing 9 Prompt - Regenerate Spatial Constraints
1prompt = f"""
2You are a spatial relationship analyzer. Given a scene description and a list of objects with their IDs, generate spatial constraints between the objects.
3
4Scene Description:
5{description}
6
7Available Objects:
8{objects_text}
9
10Last time, the following constraints were generated:
11{last_constraint_content}
12
13Edit Instructions:
14{edit_instructions}
15
16Generate spatial constraints in the following JSON format:
17[
18 {{
19 "type": "relative",
20 "relation": "right of|left of|in front of|behind|side of|on|above",
21 "source": "object_id",
22 "target": "object_id"
23 }},
24 {{
25 "type": "distance",
26 "relation": "near|far",
27 "source": "object_id",
28 "target": "object_id"
29 }},
30 {{
31 "type": "rotation",
32 "relation": "face to",
33 "source": "object_id",
34 "target": "object_id"
35 }}
36]
37
38Guidelines:
391. Only include meaningful spatial relationships that match the description
402. Use the exact object IDs from the available objects list
413. Focus on the most important spatial relationships
424. Avoid redundant relationships
435. Return only valid JSON, no additional text
446. For example, if the scene description mentions "a chair to the right of a table", the relation should be right of, with the chair as the source and the table as the target.
457. An object should either choose to be on the ground(default, no need to specify as a constraint) or on/above another object.
468. Use "on" when one object is physically resting on the surface of another object. Use "above" only when the object is floating or suspended in the air above another object, not touching it. If an object is on the ground, do not specify any "on" or "above" relation when it is the source.
47
48Generate the constraints:
49"""
Spatial Constraints Definition
Every statement extracted from the language prompt is normalized to a constraint
$c=\langle\textit{type},\;\textit{relation},\;\textit{source},\;\textit{target}\rangle,$
where source and target are object identifiers that already exist in the scene JSON. The recognized categories and their semantics are summarized in Table 3.
Type Relation Geometric Meaning
relative left of Planar ordering along the xx–axis: the source lies to the left of (negative $x$) the target’s bounding box, with a 0.1 m buffer.
right of Planar ordering along the xx–axis: the source lies to the right of (positive $x$) the target’s bounding box, with a 0.1 m buffer.
in front of Planar ordering along the yy–axis: the source lies in front of (negative $y$) the target’s bounding box, with a 0.1 m buffer.
behind Planar ordering along the yy–axis: the source lies behind (positive $y$) the target’s bounding box, with a 0.1 m buffer.
side of Planar ordering on the xx–axis: the source lies to either lateral side of the target beyond its bounding box, with a 0.1 m buffer.
distance near The horizontal gap between source and target is at most 2 m.
far The horizontal gap between source and target exceeds 8 m.
vertical on The source rests directly on the target’s top surface, with clearance under 2 mm.
above The source is positioned at least 2 m above the target, while still overlapping its footprint.
rotation face to The source’s forward-facing direction points at the target’s center within ±10°.
Table 3: Constraint primitives used by the Scene Generation Module.
Before the search phase we apply two global authoring rules:
  • Ground-only default. An object either rests on the ground or is explicitly the source of a vertical relation — never both.
  • Acyclic dependency. When generating constraints, once an object has appeared as a target it cannot later re-appear as a source, yielding a partial order that the solver exploits.
All constraints are collected into a flat JSON list with duplicate and self-referential entries removed.
DFS Solver Implementation
The DFS‐based solver proceeds as follows:
  • Preprocessing. We first build a lookup table of each object’s size, initial pose, and applicable constraints, then perform a topological sort of the objects according to the acyclic dependency rule (any object that appears as a target may not later serve as a source).
  • Candidate Generation. For each object in sorted order, we collect its active constraints against already‐placed objects.
    • If it has no constraints, we sample positions in a small neighbourhood around its initial location.
    • If it has exactly one constraint, we invoke the relation‐specific generator.
    • If it has multiple constraints, we generate candidates from the highest‐priority constraint and then filter out any that violate the others.
  • Collision and Constraint Checking. Each candidate placement is immediately tested for mesh‐level collisions and for precise satisfaction of all its constraints. Invalid placements are discarded before recursion.
  • Recursive Search and Backtracking. We place the object, recurse to the next one, and upon return remove it if we need to explore other branches. The search is depth‐first, but we track a global node counter and elapsed time: if either limit is reached, we abort further recursion and return the best complete (or partial) layout found so far.
  • Solution Selection. Whenever all objects have been placed, we compare the solution’s score (number of successfully placed items) against our current best. The highest‐scoring layout is stored and ultimately returned when the search finishes or is cut short.
Algorithm 1 DFS-Based Layout Solver
Input: Set of items $I$, spatial constraints $C$
Parameter: Timeout $T$, node limit NN
Output: Best placement $\mathrm{best\_sol}$
1: Build map of each item’s size, initial position, and rotation
2: Topologically sort II into ordered list $O$ (respecting acyclic dependencies)
[…]
4: Record $\mathrm{start\_time}\leftarrow$ now, $\mathrm{node\_cnt}\leftarrow 0$
5:procedure DFS(ii, placed)
6:if $i=|O|$ then
7:  Update $\mathrm{best\_sol}$ if current score $\mathrm{best\_score}$
8:  return
9:end if
[…]
11: Collect all constraints in $C$ with source =x=x and target $\in$ placed
12:if no such constraints then
13:  Generate candidate positions around $x$’s initial location
14:else
15:  if exactly one constraint then
16:   Generate candidate positions according to that single relation
17:  else
18:   Generate candidate positions by the primary constraint, then filter by the rest
19:  end if
20:end if
21:for each candidate in candidates do
22:  if collision detected or any constraint check fails then
23:   continue
24:  end if
25:  placed[$x$] $\leftarrow$ candidate
26:  DFS($i+1$, placed)
27:  remove $x$ from placed
28:  if timeout reached or $\mathrm{node\_cnt}>N$ then
29:   return
30:  end if
31:end for
32:return
33:end procedure
34: DFS(0, $\{\}$)
35:return $\mathrm{best\_sol}$
Details of the Scene Editing Module
After a scene is generated, users can directly edit the scene through natural language. If a user wants to adjust the spatial relationship of a specific object, they can issue a language command, which will be incorporated into the prompt as an edit instruction. GPT-o3 then uses this prompt to identify the corresponding object from the JSON file and generate new spatial constraints for it. The DFS solver treats all other objects as already placed and searches for a new position for the edited object that satisfies the updated constraints. The prompt used for this process is shown below.
Listing 10 Prompt - Scene Editing Module
1prompt = f"""
2You are a spatial relationship analyzer. Given a list of objects with their IDs and a human feedback text, extract spatial constraints.
3
4Your first task is to identify exactly one focus object from the feedback (the only object to be modified). Map this focus object to its exact ID from the Available Objects list.
5
6
7Available Objects:
8{objects_text}
9
10Human Feedback:
11{feedback}
12
13Generate spatial constraints in the following JSON format:
14[
15 {{
16 "type": "relative",
17 "relation": "right of|left of|in front of|behind|side of|on|above",
18 "source": "object_id",
19 "target": "object_id"
20 }},
21 {{
22 "type": "distance",
23 "relation": "near|far",
24 "source": "object_id",
25 "target": "object_id"
26 }},
27 {{
28 "type": "rotation",
29 "relation": "face to",
30 "source": "object_id",
31 "target": "object_id"
32 }}
33]
34
35Guidelines:
361. Identify exactly one focus object from the feedback; it must be the only "source" in all constraints. Never use the focus object as a "target".
372. Generate constraints only if they are explicitly supported by the feedback.
383. Use the exact object IDs from the Available Objects list. Ignore unknown objects.
394. Avoid redundant or contradictory relationships. If conflicts arise, choose the most specific interpretation.
405. Example: if the feedback says "Put the chair to the right of the table", use relation "right of", with the chair as the source and the table as the target.
416. Use "on" only when an object is physically resting on the surface of another object. Use "above" only when the object is suspended or floating above another object without contact. If an object is simply on the ground, do not produce "on" or "above" with it as the source.
427. Return only valid JSON, no additional text
43
44Generate the constraints:
45"""
If the user wants to replace an object, the Scene Editing Module automatically calls the 3D asset generation model to regenerate the corresponding object and replaces the original with the new one. If the user wants to add an object, the Scene Editing Module automatically calls the 3D asset generation model to generate the new object, then creates spatial constraints for it and uses the DFS solver to find a suitable placement. As for deleting objects, since all objects in the scene are manipulable mesh assets, this operation is trivial.
About the Input Type
Holodeck 2.0 is capable of generating scenes from finely detailed textual descriptions, producing results that closely match the input. However, we also recognize that in some cases, users may not have highly specific requirements and may provide only a brief, high-level description. In such scenarios, we can leverage an additional LLM to expand the user’s simple input into a more detailed scene description, which can then be used to generate the corresponding scene.
Holodeck 2.0 also supports generating a scene directly from a single image, effectively using the scene reference image as the starting point for scene generation.
Rendering Options
To achieve better rendering quality, we choose to perform the rendering in Blender. We adjust the camera among three candidate viewpoints—an overhead isometric view from the left, a frontal view, and a right‐side view—and select the one that best showcases the entire scene. The specific parameters are detailed below:
All renders use a resolution of 3840×2160 px, with 1024 samples per pixel under the Cycles engine. GPU acceleration is enabled to speed up sampling, and Blender’s built-in denoising is turned on to reduce noise while preserving fine details.
Gaming
We import the generated GLB-format scene file into Unreal Engine 5.6 for game modeling. We present additional in-game images in Figure 10 to demonstrate the richness, high quality, and interactivity of the scenes generated by Holodeck 2.0.
Qualitative Examples
In Figure 11 and Figure 12, we present 18 additional scenes generated by Holodeck 2.0, each accompanied by the fine-grained input text and the rendered scene. In Figure 13 and Figure 14, we also present 10 additional examples of scene editing, showing the original scene, the input editing instruction, and the final edited scene.
Details of Evaluation
We conduct a total of four online questionnaires; the first two focus on static scene generation in indoor and outdoor environments, respectively. Each of these two questionnaires comprises 30 distinct scenarios. For each scenario, participants answer four separate rating questions on a 10-point Likert scale (1 = worst, 10 = best), measuring:
  • Object‐level quality of Holodeck v.s. Holodeck 2.0
  • Scene‐level quality of Holodeck v.s. Holodeck 2.0
Two example items from these questionnaires are shown below:
Text: “A squat rack stands on rubber flooring, with two vertical weight plate holders attached to its sides. A flat bench is placed parallel in front of the rack at a distance of 1.2 meters. To the right of the rack, a horizontal dumbbell rack, 2 meters wide, holds dumbbells arranged from light to heavy.”
The two rendered 3D scenes generated from this text are shown below.
(a) Rendered Image of 3D Scene A
(b) Rendered Image of 3D Scene B
  1. 1.
    Please rate how well the objects generated in Scene A match the text input.
  2. 2.
    Please rate how well the overall scene in Scene A matches the text input.
  3. 3.
    Please rate how well the objects generated in Scene B match the text input.
  4. 4.
    Please rate how well the overall scene in Scene B matches the text input.
Text: “Realistic Style. A rural roadside scene includes a wooden fruit stand stocked with crates. A pickup truck is parked beside it. A folding table holds jars of jam, and a water barrel rests behind the stand on gravel ground.”
The two rendered 3D scenes generated from this text are shown below.
(a) Rendered Image of 3D Scene A
(b) Rendered Image of 3D Scene B
  1. 1.
    Please rate how well the objects generated in Scene A match the text input.
  2. 2.
    Please rate how well the overall scene in Scene A matches the text input.
  3. 3.
    Please rate how well the objects generated in Scene B match the text input.
  4. 4.
    Please rate how well the overall scene in Scene B matches the text input.
The third questionnaire focuses on the ablation study and contains a total of 60 questions. In each question, participants are asked to evaluate the similarity between two rendered images(one for Holodeck 2.0, one for using only the initial layout inferred by GPT-o3) and a given text, and to answer a single-choice question. The options for each question are: “Prefer A,” “Prefer B,” and “Tie.”
An example related to the ablation study is shown below.
Text: “Realistic Style. A winding cobblestone street intersects a small plaza dotted with stone planters. Wooden benches circle a low stone fountain. A row of trimmed hedges lines one side, while scattered fallen leaves drift across the ground. A lone bicycle leans against a lamppost.”
The two rendered 3D scenes generated from this text are shown below.
(a) Rendered Image of 3D Scene A
(b) Rendered Image of 3D Scene B
Which scene do you prefer?
  1. 1.
    Prefer A
  2. 2.
    Prefer B
  3. 3.
    Tie
The fourth questionnaire focuses on the comparison analysis of scene editing tasks. We perform 10 representative placement editing tests for both BlenderAlchemy and Holodeck 2.0. Annotators are then required to choose their preferred version between the two edited scenes. An example related to the comparison analysis is shown below.
The original scene is shown below:
Rendered image of the original scene
Please determine which of the following edited 3D scenes best matches the editing instruction: “Move the robot to the right of the crane.”
(a) Scene A
(b) Scene B
Which scene do you prefer?
  1. 1.
    Prefer A
  2. 2.
    Prefer B
Figure 10: In-game interface of the 3D scene generated by Holodeck 2.0
Figure 11: Scenes generated by Holodeck 2.0 and the corresponding fine-grained input text
Figure 12: Scenes generated by Holodeck 2.0 and the corresponding fine-grained input text
Figure 13: Edited scenes produced by the Scene Editing Module using the instructions
Figure 14: Edited scenes produced by the Scene Editing Module using the instructions